1. Consider a single workstation comprising one machine. (i) What will be the production rate of this workstation, assuming all parts processed on it take (ii) Assume that the demand is 10/day, where the day has 8 working hours, what will be the (ii) What will be the utilization of the workstation given the demand in (ii)? 20 min? throughput? (iv) If a 2nd identical machine is added to the workstation, what will be the capacity of the workstation for a 8 hour day, assuming the system is not demand constrained? Assuming that there are 100 jobs to be processed on this workstation during one shift of 8 hrs, how many machines should be included in the workstation to meet the demand? What will be the utilization for this workstation? (v)
Answer #1

1) Production rate per hour = Time available/Processing time = 60/20 = 3 units per hour

2) Throughput rate per hour = Daily demand/Time available per day = 10/8 = 1.25 units per hour

3) Utilization = Throughput rate/Production rate = 1.25/3 = 0.42

4) Capacity per day = Number of machines*Total time available in minutes per day/Processing time = 2*8*60/20 = 48 units per day